A dockworker loading crates on a ship finds that a 20-kg crate, initially at rest on a horizontal surface, requires a 73-N horizontal force to set it in motion. However, after the crate is in motion, a horizontal force of 51 N is required to keep it moving with a constant speed. Find the coefficients of static and kinetic friction between crate and floor.
Answer #1

As before it starts moving, the friction force will be static and Fs= uMg. Hence u(static)= Fs/uMg. i.e. 73/(20×9.81). i.e. 0.37.

After it starts moving kinectic friction will be in picture hence u(kinetic)= 51/196.2= 0.26. Fk/Mg.